# FRoundation: Are Foundation Models Ready for Face Recognition?

### v2

@@ 5.4 Bias evaluations foundation models FR
We evaluated the baseline model, CLIP and DINOv2 on the Racial Face in the Wild (RFW) dataset [44] to assess the model’s bias and their performance across different demographic groups. The RFW dataset contains four testing subsets corresponding to Caucasian, Asian, Indian, and African demographic groups. Following [45, 44], we reported the results as verification accuracies in (%) on each subset and as average accuracies to evaluate general recognition performance. To evaluate the bias, we reported the standard deviation (STD) between all subsets and the skewed error ratio (SER), which is given by , is the demographic group, as reported in [45, 43]. A higher STD value indicates more bias across demographic groups and vice versa. For SER, the model that achieved a value closer to 1 is less biased. Table 4 presents the evaluation results on RFW. As baseline models, we reported the results for ViT-S and ViT-L (noted as Baseline) trained from scratch on CAISA-WebFace, MS1MV2, and WebFace4M. We also reported the results of DINOv2 and CLIP without finetuning and for the case where the models are finetuned on CAISA-WebFace, MS1MV2 or WebFace4M. One can observe the following from the reported results in Table 4:
Finetuing DINOv2 and CLIP improved the general recognition performance on all demographics, in comparison to the case where pretrained DINOv2 and CLIP are used without finetuning. These results are complementary to the ones reported in the previous section.
In general, all models achieved their best recognition performances on the Caucasian subset and, in most of the settings. the lowest recognition performances on the African subset.
Using CASIA-WebFace (0.5m images) for training/finetuning, the finetuned DINOv2 and CLIP achieved higher recognition performances than the model trained from scratch. However, DINOv2 and CLIP are slightly more biased than baseline models. For example, the STD achieved by baseline (Vit-S) was 4.38, which is slightly lower than the 4.49 and 4.43 STD achieved by finetuned DINOv2 (Vit-S) and CLIP (Vit-B).
The models trained from scratch on large-scale datasets, MS1MV2 (5.8M images) and WebFace4M (4m images) archived higher average recognition performances (average) and lower standard deviation (in most of the settings) than finetuned CLIP and DINOv2. For example, using Vit-L architecture trained/finetuned on WebFace4M, the baseline achieved an average accuracy of 93.80%, in comparison to 93.43% and 91.57% achieved by DINOv2 and CLIP, respectively. Also, in this case, the baseline model achieved a lower STD (1.74) than DINOv2 (1.78) and CLIP (2.11).
In general, fined-tuned DINOv2 achieved higher average accuracies and lower STD than finetuned CLIP.
Fientuning larger model architectures of pretrained DINOv2 (Vit-L) or CLIP (Vit-L) achieved higher recognition accuracies and lower STD than finetuning smaller architectures, DINOv2 (Vit-S) or CLIP (Vit-B).
@@
| Method | Backbone | Train data | African | Asian | Caucasian | Indian | Avg. | STD | SER |
| Baseline | Vit-S | CASIA-WebFace | 75.25 | 75.68 | 84.75 | 78.58 | 78.57 | 4.38 | 1.62 |
| MS1MV2 | 96.65 | 96.32 | 98.63 | 96.68 | 97.07 | 1.05 | 2.68 | ||
| WebFace4M | 94.40 | 94.12 | 97.73 | 95.02 | 95.32 | 1.65 | 2.59 | ||
| Vit-L | CASIA-WebFace | 71.87 | 73.47 | 81.08 | 74.23 | 75.16 | 4.06 | 1.48 | |
| MS1MV2 | 97.32 | 96.48 | 98.45 | 97.25 | 97.38 | 0.81 | 2.27 | ||
| WebFace4M | 93.37 | 92.25 | 96.33 | 93.27 | 93.80 | 1.75 | 2.11 | ||
| DINOv2 | Vit-S | - | 54.77 | 61.13 | 65.00 | 60.42 | 60.33 | 4.21 | 1.29 |
| CASIA-WebFace | 76.15 | 76.90 | 85.98 | 80.65 | 79.92 | 4.49 | 1.70 | ||
| MS1MV2 | 83.43 | 83.77 | 91.18 | 87.05 | 86.36 | 3.60 | 1.87 | ||
| WebFace4M | 80.83 | 82.90 | 88.50 | 84.77 | 84.25 | 3.25 | 1.66 | ||
| Vit-L | - | 58.46 | 64.20 | 67.47 | 60.93 | 62.77 | 3.91 | 1.27 | |
| CASIA-WebFace | 85.97 | 84.00 | 93.15 | 86.65 | 87.44 | 3.96 | 2.33 | ||
| MS1MV2 | 93.75 | 93.40 | 97.08 | 93.92 | 94.54 | 1.70 | 2.26 | ||
| WebFace4M | 92.85 | 91.75 | 95.97 | 93.40 | 93.49 | 1.78 | 2.04 | ||
| CLIP | Vit-B | - | 70.75 | 69.73 | 79.32 | 68.98 | 72.19 | 4.80 | 1.49 |
| CASIA-WebFace | 80.13 | 80.53 | 89.18 | 80.30 | 82.54 | 4.43 | 1.83 | ||
| MS1MV2 | 85.60 | 86.30 | 91.82 | 87.40 | 87.78 | 2.79 | 1.76 | ||
| WebFace4M | 84.43 | 84.62 | 90.80 | 85.97 | 86.45 | 2.97 | 1.69 | ||
| Vit-L | - | 74.03 | 72.15 | 82.60 | 73.15 | 75.48 | 4.80 | 1.60 | |
| CASIA-WebFace | 84.65 | 84.47 | 92.60 | 85.02 | 86.69 | 3.94 | 2.09 | ||
| MS1MV2 | 90.63 | 90.77 | 95.03 | 91.92 | 92.09 | 2.04 | 1.88 | ||
| WebFace4M | 90.40 | 90.28 | 94.73 | 90.90 | 91.57 | 2.11 | 1.84 |
@@ 6 Conclusion
This paper was the first to propose and investigate the use of foundation models for the task of face recognition. We additionally propose the adaption of the foundation models to this specific task under various levels of data availability. Our experiments on multiple foundation models, training datasets, and a wide range set of evaluation benchmarks led to interesting conclusions summarized in the following. The studied pretrained foundation models, used as feature extractors without fine-tuning, demonstrated relatively acceptable (far from random) accuracy on less challenging face verification benchmarks like LFW. However, they under-performed on more challenging benchmarks such as AgeDB30, CALFW, IJB-B, and IJB-C. Our adaption of foundation models to face recognition showed that fine-tuning these models on even small subsets of CASIA-WebFace, e.g. 1k identities, significantly boosts their verification accuracy across benchmarks, outperforming models trained from scratch on similar data subsets. Additionally, increasing the dataset size (dataset width) or the number of images per identity (dataset depth) enhances performance, demonstrating the scalability of foundation models with diverse data availability. The results additionally showed that, with a limited training dataset like CASIA-WebFace, fine-tuning pretrained foundation models outperforms training models from scratch in recognition accuracy. However, when large datasets (MS1MV2 or WebFace4M) are available, training from scratch yields competitive performance, though it incurs a significantly higher training computational cost compared to fine-tuning. Our bias evaluation results indicate that fine-tuning foundation models enhance recognition performance across demographics but introduces slightly more bias compared to baseline models trained from scratch, especially on smaller fine-tuning datasets. Models trained from scratch on larger datasets (MS1MV2 and WebFace4M) achieved superior recognition accuracy and exhibited lower bias. The outcomes of this work set the stage for broader adoption of foundation models as a basis for biometric recognition, particularly in scenarios with limited data availability, being aware of the technical and legal constraints on biometric data collection and management.

### v3

@@ 5.5 Bias evaluations foundation models FR
| Method | Backbone | Train data | African | Asian | Caucasian | Indian | Avg. | STD | SER | Avg. | IJB-C |
| SB | 1e-4 | ||||||||||
| Baseline | ViT-S | CASIA-WebFace | 75.25 | 75.68 | 84.75 | 78.58 | 78.57 | 4.38 | 1.62 | 88.56 | 36.77 |
| MS1MV2 | 96.65 | 96.32 | 98.63 | 96.68 | 97.07 | 1.05 | 2.68 | 96.18 | 94.51 | ||
| WebFace4M | 94.40 | 94.12 | 97.73 | 95.02 | 95.32 | 1.65 | 2.59 | 96.17 | 95.62 | ||
| ViT-L | CASIA-WebFace | 71.87 | 73.47 | 81.08 | 74.23 | 75.16 | 4.06 | 1.48 | 87.39 | 62.51 | |
| MS1MV2 | 97.32 | 96.48 | 98.45 | 97.25 | 97.38 | 0.81 | 2.27 | 95.88 | 92.49 | ||
| WebFace4M | 93.37 | 92.25 | 96.33 | 93.27 | 93.80 | 1.75 | 2.11 | 95.65 | 94.79 | ||
| DINOv2 | ViT-S | - | 54.77 | 61.13 | 65.00 | 60.42 | 60.33 | 4.21 | 1.29 | 64.70 | 7.44 |
| CASIA-WebFace | 76.15 | 76.90 | 85.98 | 80.65 | 79.92 | 4.49 | 1.70 | 90.94 | 69.50 | ||
| MS1MV2 | 83.43 | 83.77 | 91.18 | 87.05 | 86.36 | 3.60 | 1.87 | 91.98 | 85.25 | ||
| WebFace4M | 80.83 | 82.90 | 88.50 | 84.77 | 84.25 | 3.25 | 1.66 | 91.46 | 85.13 | ||
| ViT-L | - | 58.46 | 64.20 | 67.47 | 60.93 | 62.77 | 3.91 | 1.27 | 66.49 | 7.84 | |
| CASIA-WebFace | 85.97 | 84.00 | 93.15 | 86.65 | 87.44 | 3.96 | 2.33 | 94.28 | 84.45 | ||
| MS1MV2 | 93.75 | 93.40 | 97.08 | 93.92 | 94.54 | 1.70 | 2.26 | 96.01 | 93.94 | ||
| WebFace4M | 92.85 | 91.75 | 95.97 | 93.40 | 93.49 | 1.78 | 2.04 | 96.03 | 94.79 | ||
| CLIP | ViT-B | - | 70.75 | 69.73 | 79.32 | 68.98 | 72.19 | 4.80 | 1.49 | 82.64 | 32.40 |
| CASIA-WebFace | 80.13 | 80.53 | 89.18 | 80.30 | 82.54 | 4.43 | 1.83 | 92.13 | 67.45 | ||
| MS1MV2 | 85.60 | 86.30 | 91.82 | 87.40 | 87.78 | 2.79 | 1.76 | 93.75 | 86.31 | ||
| WebFace4M | 84.43 | 84.62 | 90.80 | 85.97 | 86.45 | 2.97 | 1.69 | 93.11 | 85.63 | ||
| ViT-L | - | 74.03 | 72.15 | 82.60 | 73.15 | 75.48 | 4.80 | 1.60 | 86.44 | 44.69 | |
| CASIA-WebFace | 84.65 | 84.47 | 92.60 | 85.02 | 86.69 | 3.94 | 2.09 | 94.26 | 83.12 | ||
| MS1MV2 | 90.63 | 90.77 | 95.03 | 91.92 | 92.09 | 2.04 | 1.88 | 95.59 | 93.66 | ||
| WebFace4M | 90.40 | 90.28 | 94.73 | 90.90 | 91.57 | 2.11 | 1.84 | 95.59 | 93.40 |
We evaluated the baseline model, CLIP, and DINOv2 on the Racial Face in the Wild (RFW) dataset [62] to assess the models’ bias and their performance across different demographic groups. The RFW dataset contains four testing subsets corresponding to:Caucasian, Asian, Indian, and African demographic groups. Following [63, 62], we reported the results as verification accuracies in (%) on each subset and as average accuracies to evaluate general recognition performance. To evaluate the bias, we reported the standard deviation (STD) between all subsets and the skewed error ratio (SER), which is given by
, where represents the demographic group, as reported in [63, 61]. A higher STD value indicates more bias across demographic groups and vice versa. For SER, the model that achieved a value closer to 1 is less biased. Table 6 presents the evaluation results on RFW. As baseline models, we report the results for ViT-S and ViT-L (noted as baseline) trained from scratch on CASIA-WebFace, MS1MV2, and WebFace4M. We also reported the results of DINOv2 and CLIP without fine-tuning and for the case where the models are fine-tuned on CASIA-WebFace, MS1MV2, or WebFace4M. One can observe the following from the reported results in Table 6:
Fine-tuning DINOv2 and CLIP improved the general recognition performance on all demographics, in comparison to the case where pre-trained DINOv2 and CLIP are used without fine-tuning. These results are complementary to the ones reported in the previous section.
In general, the presented models achieved unequal verification performances on different subsets of RFW, where all models achieved their best verification performances on the Caucasian subset. This can be observed from the high STD values (far from optimal zero) and SER values that are far from 1. These observations are aligned with the previous works [63, 61, 44, 25, 62], reporting bias in FR when trained/finetuned on unbalanced datasets. Potential causes of this bias [67] link to the unbalanced training datasets, model’s sensitivity to skin color [39] and/or hairstyles [49].The racial distribution of the current FR datasets [66, 75, 27] used in the literature is not balanced with a majority of identities being Caucasians [63, 61, 62]. For example, as reported in [62], CASIA-WebFace [66] and MS-Celeb-1M [27], contain 84.5% and 76.3% Caucasians, 2.6% and 6.6 Asian, 1.6% and 2.6% Indian and 11.3% and 14.5% African, respectively. This is also true for the widely used WebFace260M [75] and its subsets, such as WebFace42M and WebFace4M, where the authors reported that the majority of identities are Caucasian. The previous works [63, 61, 44, 25, 62] reported that bias in these datasets would reflect in the FR algorithms, leading to higher verification accuracies on Caucasians compared to other ethnicities.
Using CASIA-WebFace (M images) for training/fine-tuning, the fine-tuned DINOv2 and CLIP achieved higher recognition performances than the baseline model trained from scratch. However, DINOv2 and CLIP are slightly more biased than baseline models. For example, the STD achieved by the baseline (ViT-S) was , which is slightly lower than the and STD achieved by fine-tuned DINOv2 (ViT-S) and CLIP (ViT-B).
The models trained from scratch on large-scale datasets, MS1MV2 (M images) and WebFace4M (M images) achieved higher average recognition performances and lower standard deviation (in most of the settings) than fine-tuned CLIP and DINOv2. For example, using the ViT-L architecture trained/fine-tuned on WebFace4M, the baseline achieved an average accuracy of %, in comparison to % and % achieved by DINOv2 and CLIP, respectively. Also, in this case, the baseline model achieved a lower STD () than DINOv2 () and CLIP ().
In general, fine-tuned DINOv2 achieved higher average accuracies and lower STD than fine-tuned CLIP.
Fine-tuning larger model architectures of pre-trained DINOv2 (ViT-L) or CLIP (ViT-L) achieved higher recognition accuracies and lower STD than fine-tuning smaller architectures: DINOv2 (ViT-S) or CLIP (ViT-B).
A higher STD and an SER further away from the optimal value of 1 indicate that there is more room for improvement in achieving uniform performance across demographics. SER values are more sensitive to slight changes in the accuracy when the error margins are small (see Equation 4). This causes models trained or fine-tuned on MS1MV2 and WebFace4M to generally have higher SER values compared to those trained or fine-tuned on CASIA-WebFace, while typically maintaining a lower STD value. This is primarily due to the higher accuracy range of the models trained on the MS1MV2 and WebFace4M datasets (resulting in lower errors), which makes the SER more sensitive to small performance differences across demographic groups.
When comparing the average verification performance on small benchmarks and IJB-C between the same model trained on WebFace4M and MS1MV2, we observe that both have comparable results in most settings. However, this observation does not hold for the performance on RFW, as the models trained on MS1MV2 outperformed those trained on WebFace4M in all cases and sometimes by a significant margin. For example, the baseline ViT large shows a performance gap of approximately 4%. This can be attributed [62] to the identity overlap between RFW and MS1MV2, a cleaned version derived from MS-Celeb-1M [27].
@@ 6 Conclusion
This paper was the first to propose and investigate the use of foundation models for the task of FR. We additionally propose the adaptation of the foundation models to this specific task under various levels of data availability. Our experiments on multiple foundation models, training datasets, and a wide range of evaluation benchmarks led to interesting conclusions, which are summarized as follows. The studied pre-trained foundation models used as feature extractors without fine-tuning, demonstrated relatively acceptable (far from random) accuracy on less challenging face verification benchmarks like LFW. However, they underperformed on more challenging benchmarks such as AgeDB30, CALFW, IJB-B, and IJB-C. Our adaptation of foundation models for FR showed that fine-tuning these models on even small subsets of CASIA-WebFace, e.g. K identities, significantly boosts their verification accuracy across benchmarks, outperforming models trained from scratch on similar data subsets. Additionally, increasing the dataset size (dataset width) or the number of images per identity (dataset depth) enhances performance, demonstrating the scalability of foundation models with diverse data availability. The results additionally showed that, with a limited training dataset like CASIA-WebFace, fine-tuning pre-trained foundation models outperforms training models from scratch in recognition accuracy. However, when large datasets (MS1MV2 or WebFace4M) are available, training from scratch yields competitive performance, though it incurs a significantly higher training computational cost compared to fine-tuning. This highlights the importance of selecting an appropriate training strategy according to the size of the dataset. Our bias evaluation results indicate that fine-tuning foundation models enhance recognition performance across demographics but introduces slightly more bias compared to baseline models trained from scratch, especially on smaller fine-tuning datasets. Models trained from scratch on larger datasets (MS1MV2 and WebFace4M) achieved superior recognition accuracy and exhibited lower bias. All models achieved their best verification performances on the Caucasian subset, which may be attributed to factors such as unbalanced training datasets, sensitivity to skin color, and hairstyles that are overrepresented or underrepresented in the training datasets. Finally, the effectiveness of synthetic data in improving the FR performance of foundation models is demonstrated, as it outperforms both pre-trained and baseline models. This also encourages further exploration of additional synthetic datasets. The outcomes of this work set the stage for broader adoption of foundation models as a basis for biometric recognition, particularly in scenarios with limited data availability, being aware of the technical and legal constraints on biometric data collection and management.
Acknowledgments
This research work has been funded by the German Federal Ministry of Education and Research and the Hessen State Ministry for Higher Education, Research and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE.